# Patterns of X-ray and -ray Flares: from Blazar to Maganetar and Sun

### v1

@@ 1 Introduction
Blazars are a distinct subclass of AGNs, characterized by their relativistic jets pointing directly toward the Earth (e.g., Ackermann etĀ al., 2015). They emit radiation across entire electromagnetic spectrum. Blazars typically exhibit strong variability, which results in prominent time-domain characteristics in multi-wavelength observations. The long-term variability of blazars has been studied through Gaussian Process method (GP; Zhang etĀ al., 2021, 2022, 2023). The main findings indicate that such variability follows the damped random walk (DRW) model, with the characteristic timescales associated with the thermal instability timescale of the accretion disk (Zhang etĀ al., 2022, 2023). As prominent blazars, both 3C 454.3 and 3C 279 exhibit -ray flaring events whose light curves were suggested to follow the stochastically driven damped simple harmonic oscillator (SHO) model in the overdamped regime (Zhang etĀ al., 2025b). This variability behavior is typically attributed to internal processes within the innermost regions of their relativistic jets (e.g., Ackermann etĀ al., 2016; Shukla etĀ al., 2018).
@@
Magnetars are a type of neutron stars with extremely strong magnetic fields ( G; see Kaspi & Beloborodov (2017) for a review), whose intense magnetic activity can trigger multi-band radiation bursts, e.g., XRBs, radio bursts (may include fast radio bursts, FRBs), and -ray bursts (Turolla etĀ al., 2015). Studies suggest that the energy release from magnetar bursts could be linked to direct magnetic reconnection in the magnetosphere, driven by the intense magnetic field (e.g., Beloborodov, 2013; Mereghetti etĀ al., 2015). Alternatively, these flares might be triggered by shear waves released during crustal rupture, which transmit energy to the magnetosphere through Alfvn waves, initiating secondary magnetic reconnection (e.g., Li & Beloborodov, 2015; Yuan etĀ al., 2020). SGR J1935+2154 is a Galactic magnetar, of which hundreds of XRBs were recorded by different detectors, e.g., Insight-HXMT (Li etĀ al., 2021), Neutron Star Interior Composition Explorer (NICER; Yuan etĀ al., 2020). A non-thermal XRB of SGR J1935+2154 on 2020 April 28th has been discovered to be accompanied by emission of the FRB 200428 (CHIME/FRB Collaboration etĀ al., 2020; Li etĀ al., 2021), suggesting the non-thermal XRB and FRB 200428 share the same physical origin in an explosive event from SGR J1935+2154. For XRBs in SGR J1935+2154, Tang etĀ al. (2024) used GP method to characterize their variability. They found that both normal XRBs and FRB-associated HXMT XRB generally show variability consistent with the SHO patern, the derived timescales can put constraints on the production mechanisms of XRBs.
@@
Solar flares are explosive phenomena in the solar atmosphere, which can emit electromagnetic emission extending to -rays within a short time. Extensive research has been conducted on particle acceleration, radiation mechanisms, and emission regions in solar flares (e.g., Holman etĀ al., 2011; Chen etĀ al., 2020; Share etĀ al., 2024). It is generally agreed that energy release is usually attributed to magnetic reconnection occurring on the current sheet (e.g., Chen, 2011; Benz, 2017; Share etĀ al., 2018; Chen etĀ al., 2020). Research indicates that the light curves of hard X-rays ( keV) display fast pulsating structures with periods ranging from milliseconds to seconds, closely related to the acceleration of high-energy electrons induced by magnetic reconnection, with non-thermal bremsstrahlung as the primary radiation mechanism (e.g., Holman etĀ al., 2011; Kontar etĀ al., 2019; Krucker etĀ al., 2020).
@@ 2 Method and data processing
The Reuven Ramaty High Energy Solar Spectroscopic Imager (RHESSI) is a NASA Small Explorer Mission Lin etĀ al. (2002). It is capable of performing both imaging and spectroscopic analyses across a broad energy spectrum, ranging from soft X-rays to -rays (3 keV to 17 MeV). The highest energy resolution is as fine as 1 keV. In this study, we utilize the SSWIDL software to extract RHESSI hard X-ray lightcurves, with the temporal resolution of flare light curves set at 4 seconds.
We perform a detailed analysis of the flare light curves using the GP method (Aigrain & Foreman-Mackey, 2023). The mean function of the GP is chosen to be a constant equal to the average count rate of the light curve for each flare. The covariance (kernel) function of the GP therefore serves to represent all variability characteristics of the flares, with its selection being a critical issue in GP. Following empirical practice, we opt for three kernel functions to fit the light curves, i.e., DRW, SHO, and Matrn.
DRW is typically used to capture random fluctuations that gradually decay, with its PSD performing a broken power-law form (e.g., MacLeod etĀ al., 2010; Burke etĀ al., 2021). SHO is most commonly used to describe periodic behaviors in time series, corresponding to the underdamped regime (with the quality factor of the oscillator ), where its PSD shows a peak (e.g., Yang etĀ al., 2021; Zhang etĀ al., 2021). In the overdamped regime (), the PSD is more complex, with the low-frequency part matching the DRW PSD, where the spectral index transitions from -2 to 0 111The change of spectrum index mentioned in this paper refers to the change from high frequencies to low frequencies. (characteristic timescale denoted as ). At high frequencies, the spectral index drops to -4 (the timescale corresponding to the transition from spectral index -4 to -2). In the critically damped regime (), the PSD tends to exhibit a direct transition from a spectral index of -4 to 0 (Kasliwal etĀ al., 2017; Moreno etĀ al., 2019). Different regimes may describe distinct characteristics of the data. The Matrn kernel can capture data structures with large fluctuations, and its PSD features a direct transition from a spectral index of -4 to 0 (Rasmussen & Williams, 2006).
For Matrn, the function is as follow:
which is derived by setting the smoothness parameter (half-integer) of the Matrn class kernels to 3/2. and represent the characteristic timescale and the amplitude respectively. The details of DRW and SHO models can refer to our previous works (Zhang etĀ al., 2021, 2022, 2023). Here, we supplement two timescale (broken timescale of PSD) formulas for the SHO model in overdamped modeļ¼
@@
we use a tool of Celerite (Foreman-Mackey etĀ al., 2017) to perform GP fitting time series. Markov chain Monte Carlo (MCMC) technology is used when implementing the light curve fitting. We implement it through the package emcee 222https://github.com/dfm/emcee (Foreman-Mackey etĀ al., 2013), taking the last 30,000 steps out of 50,000 total MCMC steps as effective samples. Actually, before running MCMC, we choose the initial parameter values through calculating the maximum likelihood. Then, we can construct the final PSD by taking the median of the 30,000 PSD values corresponding to each frequency, and plot 30,000 groups of parameters as a distribution, i.e., a posterior distribution of the model parameters.
@@
Three bright solar hard X-ray flares, occurring on 23 July 2002 (flare 1), 3 November 2003 (flare 2) and 25 February 2014 (flare 3) in the 50-100 keV and 100-300 keV energies, are analyzed, with data obtained from the RHESSI satellite. The specific time periods for the flares are listed in Table 1. DRW, SHO and Matrn models are chosen to describe these light curves. Considering the fit between the observed data and modeled light curve, the distribution of standard residuals, as well as the autocorrelation functions (ACF) of both the standard residuals and their squares (the goodness-of-fit criteria described in detail in Zhang etĀ al. (2022)), we observe no significant differences between the three models for any of the flares. Further evaluation using the corrected Akaike information criterion (), with values listed in TableĀ 1, shows that the SHO model outperforms the DRW and Matrn models with for all flares across both energy bands.
We present the SHO fitting plots for the solar flares in the 50-100 keV and 100-300 keV energy bands, as shown in FigureĀ 1 and FigureĀ 2. From left to right, the panels display the results for flare 1, flare 2 and flare 3 respectively. Although the residual distribution for the flare 3 does not fit a normal distribution well, the concentration of residuals around zero and the lack of trends or patterns suggest that the SHO model is still considered valid.
The corresponding posterior probability density distributions of the SHO parameters are shown in FigureĀ 3. The SHO parameters for all flares are well constrained, with the specific values provided in TableĀ 1.
Now we focus on the PSDs shown in FigureĀ 4. Flare 1 and flare 3 have PSD indexes changing from -4 to -2 and then to 0. That is, the timescales of and are constrained. The two timescales derived from flare 3 in 100-300 keV are in close proximity to each other. However, the PSD indexes of flare 2 in two energy bands evolve from -4 to -2, with only constrained. The spectral indexes of PSDs and the characteristic timescales of these solar flares are listed in the TableĀ 2. and of the solar flares are in the range of 4-7 s and 16-53 s, respectively. The relationship between these characteristic timescales and the broken frequency of the PSD is .
@@
flare 3 | 2014-02-25 00:43:22-00:50:02 | 1235 | 1414 | 1483 |
Note. ā (1) source name, (2) time period of the flare, (3)(4)(5) posterior parameters of modeling light curves with SHO model, (6)(7)(8) values of SHO, DRW and Matrn model. The uncertainties of model parameters represent confidence intervals. The upper part shows the results for 50-100 keV, and the lower part for 100-300 keV.
@@ 3.2 Comparison of flares from blazars, magnetars and the Sun
Considering that the exact mode of the SHO model is determined by the quality oscillation factor , we establish a two-dimensional relationship between the peak luminosity-to-mass ratio of the flares and the parameter derived from the SHO model fits, as shown in FigureĀ 5. For solar flares (the green part in FigureĀ Ā 5), the peak luminosity estimated as erg/s is used. The values of flare 1 and flare 2 are generally near or below 0.3, while flare 3 appears to have values significantly larger than 0.3 and develops to larger values (near or larger than 0.5).
@@
There is a unique XRB in SGR 1935+2154 that is associated with FRB 200428. Its high-energy (HE) data from Insight-HXMT show the best correspondence with the FRB radio data. The value of this XRB in HE is approximately 0.3, which is comparable to that of solar X-ray flares 1 2, yet its peak luminosity-to-mass ratio exceeds that of solar flares by roughly 8-10 orders of magnitude (see FigureĀ 5). The normal XRBs in SGR 1935+2154 possess values spanning a wide range (Tang etĀ al., 2024), without displaying in the figure.
For -ray blazar flares (the blue part), there is a difference of 10-11 orders of magnitude for peak luminosity-to-mass ratio compared to solar X-ray flares. They have values generally larger than 0.3, which is similar to that of solar flare 3. We add the results of -ray long-term variability of 3C 454.3 and 3C 279 (the baby blue part) as a comparison of the blazar flares. Both their values () and the peak luminosity-to-mass ratio are lower than those of the flares. Blazar PG 1553+113 and non-blazar PKS 0521-36 (the orange arrows) exhibit long-term -ray variability with a quasi-periodic oscillation (QPO) signal, and their values deviate significantly from the others shown in FigureĀ 5. It should be noted that the long-term -ray variability of PG 1553+113 is not dominated by this periodic component, but rather by a component described by the Matrn model (Zhang etĀ al., 2021, 2025a).
@@
Additionally, we provide a summary of the PSD characteristics and variability timescales for the analyzed flares in TableĀ 2. The FRB-associated HXMT-HE XRB, like solar flares, possesses a PSD structure with an index changing from -4 to -2, but the timescale of the HXMT-HE XRB is significantly smaller. As a result of data constraints, both the FRB-associated HXMT-HE XRB and solar flare 2 are missing a PSD structure with an index changing from -2 to 0. The blazar -ray flares have the longest observed characteristic timescales, with PSD indexes changing from -4 to 0.
We further simulate the PSDs of SHO for different values and show them in FigureĀ 6. At low values of , the PSD displays two distinct breaks. With an increase in , the two bending points gradually approach each other, merging into almost one break at =0.5. When , the peak feature (QPO signal) becomes visible. This can serve as a reference for the results we have fitted.

### v2

@@ 1 Introduction
Blazars are a distinct subclass of AGNs, characterized by their relativistic jets pointing directly toward the Earth (e.g., 2015ApJ...810...14A). They emit radiation across entire electromagnetic spectrum. Blazars typically exhibit strong variability, which results in prominent time-domain characteristics in multi-wavelength observations. To study such variability, the Gaussian Process (GP; 2021ApJ...919...58Z; 2022ApJ...930..157Z; 2023ApJ...944..103Z) method has been increasingly applied in recent years. GP is a powerful data analysis tool based on the theory of probability statistics and Bayesian analysis techniques, providing a flexible, non-parametric framework for modeling time series data. In GP, the kernel function encodes assumptions about the correlation structure, allowing us to capture both smooth trends and stochastic fluctuations. Results from the analyzed studies, obtained through GP modeling, indicate that the long-term variability of Blazars follows the damped random walk (DRW) model, with the characteristic timescales associated with the thermal instability timescale of the accretion disk (2022ApJ...930..157Z; 2023ApJ...944..103Z). As prominent blazars, both 3C 454.3 and 3C 279 exhibit -ray flaring events whose light curves were suggested to follow the stochastically driven damped simple harmonic oscillator (SHO) model in the overdamped regime (2025MNRAS.540.3790Z). This variability behavior is typically attributed to internal processes within the innermost regions of their relativistic jets (e.g., 2016ApJ...824L..20A; 2018ApJ...854L..26S).
@@
Magnetars are a type of neutron stars with extremely strong magnetic fields ( G; see 2017ARA&A..55..261K for a review), whose intense magnetic activity can trigger multi-band radiation bursts, e.g., XRBs, radio bursts (may include fast radio bursts, FRBs), and -ray bursts (2015RPPh...78k6901T). Studies suggest that the energy release from magnetar bursts could be linked to direct magnetic reconnection in the magnetosphere, driven by the intense magnetic field (e.g., 2013ApJ...762...13B; 2015SSRv..191..315M). Alternatively, these flares might be triggered by shear waves released during crustal rupture, which transmit energy to the magnetosphere through Alfvn waves, initiating secondary magnetic reconnection (e.g., 2015ApJ...815...25L; 2020ApJ...900L..21Y). SGR J1935+2154 is a Galactic magnetar, of which hundreds of XRBs were recorded by different detectors, e.g., Insight-HXMT (2021NatAs...5..378L), Neutron Star Interior Composition Explorer (NICER; 2020ApJ...900L..21Y). A non-thermal XRB of SGR J1935+2154 on 2020 April 28th has been discovered to be accompanied by emission of the FRB 200428 (2020Natur.587...54C; 2021NatAs...5..378L), suggesting the non-thermal XRB and FRB 200428 share the same physical origin in an explosive event from SGR J1935+2154. For XRBs in SGR J1935+2154, 2024ApJ...971...26T used GP method to characterize their variability. They found that both normal XRBs and FRB-associated HXMT XRB generally show variability consistent with the SHO patern, the derived timescales can put constraints on the production mechanisms of XRBs.
@@
Solar flares are explosive phenomena in the solar atmosphere, which can emit electromagnetic emission extending to -rays within a short time. Extensive research has been conducted on particle acceleration, radiation mechanisms, and emission regions in solar flares (e.g., 2011SSRv..159..107H; 2020NatAs...4.1140C; 2024arXiv241219586S). It is generally agreed that energy release is usually attributed to magnetic reconnection occurring on the current sheet (e.g., 2011LRSP....8....1C; 2017LRSP...14....2B; 2018ApJ...869..182S; 2020NatAs...4.1140C). Research indicates that the light curves of hard X-rays ( keV) display fast pulsating structures with periods ranging from milliseconds to seconds, closely related to the acceleration of high-energy electrons induced by magnetic reconnection, with non-thermal bremsstrahlung as the primary radiation mechanism (e.g., 2011SSRv..159..107H; 2019ApJ...871..225K; 2020ApJ...894..158K).
@@ 2 Method and data processing
The Reuven Ramaty High Energy Solar Spectroscopic Imager (RHESSI) is a NASA Small Explorer Mission 2002SoPh..210....3L. It is capable of performing both imaging and spectroscopic analyses across a broad energy spectrum, ranging from soft X-rays to -rays (3 keV to 17 MeV). The highest energy resolution is as fine as 1 keV. In this study, we extract hard X-ray lightcurves using the SSWIDL software, based on the RHESSI Observing Summary Data. The summary lightcurves provide count rates in nine pre-selected energy bands with a temporal resolution of 4 s. Heuristic corrections are applied to remove the effects of attenuator and decimation state changes; these corrections are sufficient for timing and variability studies, although they do not provide absolute photon fluxes which are generally required only for detailed spectral analysis. For our analysis, we focus on high-energy bands where the effect of pile-up is less pronounced, and therefore do not explicitly consider attenuator status, which mainly affects low-energy (40 keV) observations. All light curves were generated using detectors 1, 3, 4, 5, 6, and 9, which are known for their stable performance and low noise levels. We conduct a detailed analysis of the flare light curves within the GP framework (2023ARA&A..61..329A), focusing on the near-peak phase of the solar flares, which approximately corresponds to the impulsive phase (highlighted by the gold-shaded region in Figure 1).
GP can be regarded as a distribution over functions, fully specified by a mean function and a covariance (kernel) function. Instead of fitting a fixed functional form, GP places a prior directly on the space of possible functions, and the kernel determines how data points are correlated in time. This property makes GP particularly suitable for irregularly sampled and noisy astrophysical light curves. In addition, GP is capable of characterizing fast variabilities, such as those frequently associated with quasi-periodic pulsations (QPPs) which are commonly observed during the impulsive phase of solar flares. In our modeling, the mean function of the GP is chosen to be a constant equal to the average count rate of the light curve for each flare. The kernels act as key components of the GP framework. They are mathematical models whose primary function is to quantify the similarity between data points in time, thereby governing the structure and smoothness of the resulting light-curve models. Following empirical practice, we opt for three kernel functions to fit the light curves, i.e., DRW, SHO, and Matrn which are used in our previous works (2021ApJ...919...58Z; 2022ApJ...930..157Z; 2023ApJ...944..103Z).
The DRW process can be described by the following first-order stochastic differential equation (2009ApJ...698..895K):
where represents the amplitude term, is correlation decay timescale or damping timescale, and represents a white noise process. The kernel function of the DRW is written as:
The power spectral density (PSD) of this process is performing as a broken power-law form (e.g., 2010ApJ...721.1014M; 2021Sci...373..789B):
with the spectral index transiting from -2 to 0111The change of spectrum index mentioned in this paper refers to the change from high frequencies to low frequencies.. The DRW kernel corresponds to the Ornstein-Uhlenbeck (OU) process (PhysRev.36.823) in physics. It is typically used to capture random fluctuations that gradually decay.
The form for a SHO process driven by a random (white noise) forcing is:
where is the natural frequency of the undamped oscillator, and is its quality factor of the oscillator. The SHO kernel function is relatively sophisticated, and we refer the reader to 2017AJ....154..220F for a more detailed description. The PSD of this process is
where . The SHO represents a more complex model that can be categorized into three regimes according to the value of the oscillation quality factor . When , it falls in the overdamped regime. In this regime, the PSD matches the DRW PSD at low frequencies, i.e., the spectral index transitions from -2 to 0 (characteristic timescale denoted as ), and the spectral index drops to -4 at high frequencies (where the timescale corresponds to the transition of the spectral index from -4 to -2). The two characteristic timescales can be calculated as:
When , SHO falls in the underdamped regime, where its PSD shows a peak (e.g., 2021ApJ...907..105Y; 2021ApJ...919...58Z; 2025MNRAS.537.2380Z), representing a quasi-periodic oscillation (QPO) signal. When , it falls in the critically damped regime, with the spectral index of PSD tending to transform from -4 directly to 0 (refer to Figure 7 for a more intuitive representation; 2017MNRAS.470.3027K; 2019PASP..131f3001M). The SHO model is most commonly used to describe periodic behaviors in time series. Actually, depending on the regime, it can capture distinct characteristics of the variability.
When setting the smoothness parameter (half-integer) of the Matrn class kernels to 32, we get Matrn. This process can be described by a second-order stochastic differential equation (2012Kalman):
represents white noise process, is the same as in DRW model, representing correlation decay timescale. The kernel function of Matrn is:
where represents the amplitude terms. The corresponding PSD can be written as (2006gpml.book.....R; 2012Kalman):
with its spectral index transforming from -4 to 0. Compared to the DRW (equal to the Matrn class kernel with ) model, the Matrn model produces curves that are smoother at high frequencies, and its smoothness and correlation can be adjusted by the parameter .
@@
we use a tool of Celerite (2017AJ....154..220F) to perform GP fitting time series. Markov chain Monte Carlo (MCMC) technology is used when implementing the light curve fitting. We implement it through the package emcee 222https://github.com/dfm/emcee (2013PASP..125..306F), taking the last 30,000 steps out of 50,000 total MCMC steps as effective samples. Actually, before running MCMC, we choose the initial parameter values through calculating the maximum likelihood. Then, we can construct the final PSD by taking the median of the 30,000 PSD values corresponding to each frequency, and plot 30,000 groups of parameters as a distribution, i.e., a posterior distribution of the model parameters.
@@
We select three bright solar hard X-ray flares: the X4.8-class flare on 23 July 2002 (flare 1), the X3.9-class flare on 3 November 2003 (flare 2), and the X4.9-class flare on 25 February 2014 (flare 3), as representative cases for analysis. They were selected based on the following scientific considerations: (1) Geostationary Operational Environmental Satellites (GOES) classX3 (with peak 1-8 soft X-ray flux measured by GOES exceeding ), corresponding to strong hard X-ray emission; (2) Continuous, high-quality RHESSI observations covering the entire impulsive phase, ensuring sufficient signal-to-noise for reliable analysis; (3) Pronounced and structured variability in the 50-100 keV and 100-300 keV RHESSI light curves, providing clear temporal signatures for our variability characterization framework. Furthermore, they are well-documented benchmark events in RHESSI solar flare studies (e.g., 2003ApJ...595L..97H; 2004ApJ...611L..53L; 2005A&A...434.1173G; 2023CosRe..61..265S). To provide a clearer identification of the solar flares analyzed in this study, Figure 1 present the full-duration GOES 1-8 Å soft X-ray profiles (the black solid line) together with the original RHESSI 50-100 keV (the red solid line) and 100-300 keV (the yellow solid line) hard X-ray profiles for each flare (three solar flares shown in panels (a)(b)(c) respectively), where the time intervals used in our analysis are marked (the gold shaded region). For visual comparison, a representative -ray flare light curve of blazar 3C 279 is also shown side by side (panel (d)). The specific analyzed time intervals for the solar flares are listed in Table 1.
DRW, SHO and Matrn models are chosen to describe these RHESSI hard X-ray flare light curves. From the analysis of the fit between the observed data and modeled light curve, the distribution of standard residuals, as well as the autocorrelation functions (ACF) of both the standard residuals and their squares (the goodness-of-fit criteria described in detail in 2022ApJ...930..157Z), no significant differences among the three models for any of the flares. Further evaluation using the corrected Akaike information criterion (), with values listed in Table 1, shows that the SHO model outperforms the DRW and Matrn models with for all flares across both energy bands.
We present the SHO fitting plots for the solar flares in the 50-100 keV and 100-300 keV energy bands, as shown in Figure 2 and Figure 3. From left to right, the panels display the results for flare 1, flare 2 and flare 3 respectively. Although the residual distribution for the flare 3 does not fit a normal distribution well (the bottom left panel for the third columns in Figure 2 and Figure 3), the concentration of residuals around zero and the lack of trends or patterns suggest that the SHO model is still considered valid.
The corresponding posterior probability density distributions of the SHO parameters are shown in Figure 4. The SHO parameters for all flares are well constrained, with the specific values provided in Table 1.
Now we focus on the PSDs shown in Figure 5. Flare 1 and flare 3 have PSD indexes changing from -4 to -2 and then to 0. That is, the timescales of and are constrained. The two timescales derived from flare 3 in 100-300 keV are in close proximity to each other. However, the PSD indexes of flare 2 in two energy bands evolve from -4 to -2, with only constrained. The spectral indexes of PSDs and the characteristic timescales of these solar flares are listed in the Table 2. and of the solar flares are in the range of 4-7 s and 16-53 s, respectively. The relationship between these characteristic timescales and the broken frequency of the PSD is .
@@
| flare 3 | 2014-02-25 00:43:22-00:50:02 | 1235 | 1414 | 1483 | |||
Note. — (1) source name, (2) time period of the flare, (3)(4)(5) posterior parameters of modeling light curves with SHO model, (6)(7)(8) values of SHO, DRW and Matrn model. The uncertainties of model parameters represent confidence intervals. The upper part shows the results for 50-100 keV, and the lower part for 100-300 keV.
@@ 3.2 Comparison of flares from blazars, magnetars and the Sun
Considering that the exact mode of the SHO model is determined by the quality oscillation factor , we establish a two-dimensional relationship between the parameter derived from the SHO model fits and the flare analysis timescale (i.e., the duration of the flare light curves we analyzed), as shown in Figure 6, to probe the relationship between the dynamics of the different burst events and the their local physical scale of the emission region. For solar flares (the green part in Figure 6), the values of flare 1 and flare 2 are generally near or below 0.3, while flare 3 appears to have values significantly larger than 0.3 and develops to larger values (near or larger than 0.5).
@@
There is a unique XRB in SGR 1935+2154 that is associated with FRB 200428. Its high-energy (HE) data from Insight-HXMT show the best correspondence with the FRB radio data. The value of this XRB in HE is approximately 0.3 (as shown by the red mark in Figure 6; 2024ApJ...971...26T), which is comparable to that of solar X-ray flares 1 2, yet its analysis timescale is less than that of solar flares by roughly 3 orders of magnitude. The normal XRBs in SGR 1935+2154 possess values spanning a wide range (2024ApJ...971...26T), without displaying in the figure.
For -ray blazar flares (the blue part in Figure 6; 2025MNRAS.540.3790Z), there is a difference of 3-4 orders of magnitude for compared to solar X-ray flares. They have values generally larger than 0.3, which is similar to that of solar flare 3. Their PSDs nearly exhibit a single break structure (2025MNRAS.540.3790Z).
@@
Additionally, we provide a summary of the PSD characteristics and variability timescales for the analyzed flares in Table 2. The FRB-associated HXMT-HE XRB, like solar flares, possesses a PSD structure with an index changing from -4 to -2, but the timescale of the HXMT-HE XRB is significantly smaller. As a result of data constraints, both the FRB-associated HXMT-HE XRB and solar flare 2 are missing a PSD structure with an index changing from -2 to 0. The blazar -ray flares have the longest observed characteristic timescales, with PSD indexes changing from -4 to 0.
We further simulate the PSDs of the SHO for different values, along with the corresponding simulated light curves, as shown in Figure 7, to better illustrate our results. The PSD evolution with reveals a clear sequence: two distinct breaks at low converge into one at , after which a QPO peak emerges for . We interpret this entire sequence as reflecting the interaction between two underlying timescales: one from energy injection (; high-frequency break) and the other from energy dissipation (; low-frequency break). The convergence reflects a dynamic rebalancing between the processes, guiding the system to a recovered state of equilibrium, while the eventual QPO marks the excitation of a coherent dynamical mode within the system, likely triggered by the energy injection process. Through this figure, we can not only better examine how the quality factor controls the variability features, such as the amplitude of fluctuations, but also investigate how the two characteristic timescales in the PSD evolve and reflect possible system states. In the context of our solar flare results, Flares 1 and 2 appear to occur in systems where energy is injected relatively rapidly but dissipates more slowly, whereas Flare 3 likely occurs in a near-equilibrium system, where any injected energy is immediately dissipated.